# An Analysis of Decision Problems for Relational Pattern Languages under Various Constraints

### v1

@@ 2 Preliminaries
In the proofs of the main results regarding the inclusion problem, we make use of two different automata models that have undecidable language emptiness properties. The first one is the so called nondeterministic 2-counter automaton without input (see, e.g., [Iberra1978]). The second one is a very specific universal Turing machine $U$ with 2 symbols and 15 states (see [Neary2009]). As the technical proofs are very lengthy and based on the constructions provided in other papers, the formal definitions of the models can be found in Appendix  0.B.1 and Appendix  0.B.2 respectively. For the intuitions of the proofs provided in the main body, it suffices to know that these respective models have undecidable language emptiness properties, that all terminating/accepted computations are encoded using finite binary words, and that the number of existing and not existing outgoing transitions per state are bound by a constant. This last property results in a finite number of properties of words that are not encodings of accepting computations.
3 The Membership Problem
In this section, we discuss the membership problem for relational pattern languages for the relations in $\mathcal{R}$. In addition to that, we also consider certain special cases of one-sided relations mentioned at the end of [Angluin1980] that do not immediately translate to the framework of relational pattern languages considered here. As to be seen, the membership problem remains NP-complete for all considered cases. We will begin by covering the cases arising for all relations in $\mathcal{R}$.
For the case of $r_{=}$, we recall the original result by Angluin in [Angluin1980] in which it is shown that the membership problem of non-erasing pattern languages is NP-complete. In the erasing case, the same result has been shown by Jiang et al. in [Jiang1994]. The following theorem states the main result of this section. It is the NP-completeness of the membership problem for relational pattern languages over any of the relations in $\mathcal{R}$. The rest of this section serves as its proof.
Theorem 3.1
Let $r\in\mathcal{R}$ and assume $R=\{r\}$. Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{X}(\alpha,r_{\alpha})$ for $X\in\{E,NE\}$ is NP-complete.
Indeed, the NP-containment of the membership problems for any of the relations follows by a rather similar argument as can be made for the case of the equality relation $r_{=}$ (see Appendix 0.A for more details).
Proposition 1(✂)
@@ 3 The Membership Problem
Let $r\in\mathcal{R}$ and $R=\{r\}$. Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{X}(\alpha,r_{\alpha})$ for $X\in\{E,NE\}$ is in NP.
The Non-Erasing Cases: To show NP-hardness, we begin with the non-erasing case and start with the relations for which the exact same construction as in [Angluin1980] can be applied. An overview of the original proof together with respective arguments why it works for all these other relations can be found in Appendix 0.A.
Lemma 1(✂)
Let $r\in\{r_{=},r_{|w|},r_{ssq},r_{ab},r_{perm},r_{rev},r_{*}\}$. In the case of $r=r_{hd,k}$, assume $k\in\mathbb{N}$. Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
That leaves us only with the cases of commutation. Notice that for the non-erasing cases, $r_{com^{*}}$ and $r_{com^{+}}$ result in the same valid substitutions as the empty word is excluded from valid substitutions.
Proposition 2(✂)
Let $r\in\{r_{com^{*},com^{+}}\}$. Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
Again, the full technical proof can be found in Appendix 0.A. To give an intuition, this proof also relies on a reduction from the $3\mathtt{-SAT}$ problem and has a similar nature as the proof used for Lemma˜1 which is based on the proof from [Angluin1980]. For the adapted construction to work, we only use two different kinds of factors instead of three. First, a part that handles the variable’s true and false encodings, by setting the factor in $w$ to 1#11\#1 and the factor in $\alpha$ to $u_{i}v_{i}$, for each $3\mathtt{-SAT}$ variable $X_{i}$ (resulting in only either $u_{i}$ or $v_{i}$ to be substituted to contain a #\#). For the parts handling the clauses, we use the factor $1^{10}\#1^{10}\#1^{10}$ in $w$ and the factor $z_{j,1}\ x_{j,1}\ z_{j,2}\ x_{j,2}\ z_{j,3}\ x_{j,3}\ z_{j,4}$ in $\alpha$, where each $x_{j,i}$ corresponds to some variable related to either some $u_{k}$ or $v_{k}$ and each $z_{j_{i}}$. By that, at least one variable in each clause must not contain a #\#, resulting in the reduction to go through. This concludes the non-erasing cases.
The Erasing Cases: Now, we continue with the erasing case. As in the non-erasing cases, for many relations, the existing construction (this time from [Jiang1994]) can be directly applied to show NP-hardness (see Appendix 0.A for some more details).
Lemma 2(✂)
Let $r\in\{r_{=},r_{|w|},r_{ssq},r_{ab},r_{perm},r_{rev},r_{*}\}$. In the case of $r=r_{hd,k}$, assume $k\in\mathbb{N}$. Given $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
That leaves us with ++-commutation and *-commutation. First of all, for the case of $r_{com^{+}}$, we see that the proof used to show Proposition˜2 also applies for the erasing case, as two related variables are never substituted to the empty word (see Appendix 0.A for more details).
Proposition 3(✂)
Let $R=\{r_{com^{+}}\}$. Given some word $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{NE}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
Notice that $r_{com^{*}}$, however, always allows for one of two related variables to be substituted by the empty word. Therefore, we need a slightly different construction for this to work. Based on the proof of Proposition Proposition˜2, we can change the clause encoding in $w$ to only the letter $1$ instead of $1^{10}\#1^{10}\#1^{10}$ and remove all $z$ variables for the construction to work. The basic idea is that is a variable is substituted to contain the letter #\#, it cannot be substituted to a single $1$, reversing the idea used in Proposition˜2 (See Appendix 0.A).
Proposition 4(✂)
Let $R=\{r_{com^{*}}\}$. Given some word $w\in\Sigma^{*}$ and $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, deciding whether $w\in L_{E}(\alpha,r_{\alpha})$ is NP-hard for all alphabets $\Sigma$ with $|\Sigma|\geq 2$.
This concludes the NP-hardness results for the membership problem under all relations considered and, thus, the proof of Theorem 3.1. As mentioned in the beginning of this section, we also obtain NP-hardness in the cases of restricting the constructed patterns to one-sided subsequence or one-sided star relations. We obtain the following results by proofs that are similar to the ones used for the previous results. Due to space constraints and their special nature, the discussion of the following result can be found in Appendix 0.A.1.
Theorem 3.2
Given $w\in\Sigma^{*}$ and some pattern $\alpha$ where each variable $x_{i}$ occurs at most once, and, if it does, the related variables $x^{*}$ (resp. $x^{s}$) [^1] For the discussion of restrictions to one-sided relations, we deviate from the notation used in all other parts of this paper and use a notation similar to the one introduced in the end of the work by Angluin[Angluin1980]. may occur multiple times, it is NP-complete to decide whether $w\in L_{X}(\alpha)$, where $X\in\{E,NE\}$.
@@ 4 The Inclusion Problem
In this section, we continue with the inclusion problem for relational pattern languages. We recall that the inclusion problem for non-erasing as well as erasing pattern languages over $r_{=}$ is generally undecidable for all alphabets $\Sigma$ of size $|\Sigma|\geq 2$, as shown and refined in  [Jiang1995, Freydenberger2010, BREMER201215]. As we will see in this section, for several of the considered relations in $\mathcal{R}$, we also obtain that this problem is undecidable in the non-erasing or erasing cases.
In fact, if we consider, first, the relations $r_{ssq}$ and $r_{*}$, we observe that we can simulate the exact behavior of $r_{=}$, similar to the NP-hardness proofs in the membership section. Meaning, if we take any two variables $x,y\in\operatorname{var}(\alpha)$, for some relational pattern $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, and any of the three relations above, we see that relating these variables in both direction effectively forces two variables to be substituted equally in $r_{\alpha}$ valid substitutions, i.e., if $(x,y),(y,x)\in r_{\alpha}$, then $h\in H_{r_{\alpha}}$ if and only if $h(x)=h(y)$. By the fact that this problem is undecidable for $r_{=}$, this results in the following, for which the formal proof can be found in Appendix 0.B.3 for reference.
Corollary 1
[Jiang1995, Freydenberger2010, BREMER201215] Let $\Sigma$ be any finite alphabet with $|\Sigma|\geq 2$. Let $\tilde{r}\in\{r_{=},r_{ssq},r_{*}\}$ and let $R=\{\tilde{r}\}$. Given $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is undecidable to answer whether $L_{\mathcal{X}}(\alpha,r_{\alpha})\subseteq L_{\mathcal{X}}(\beta,r_{\beta})$, for $\mathcal{X}\in\{E,NE\}$.

### v2

@@ 2 Preliminaries
  10. 10.
    $r_{\varepsilon_{=}}$ : common (non)emptiness, i.e., $(u,v)\in r_{\varepsilon_{=}}$ iff $u=\varepsilon=v$ or $u,v\in\Sigma^{+}$.
3 The Equivalence Problem
In this first main section, we cover the equivalence problem for relational pattern languages. We recall that in the erasing case, this problem is one of the main open problems for the $r_{=}$ relation. We obtain a series of distinct results.
For the non-erasing case, we give a characterization of a set of relations under which equivalence is characterized by equality of patterns and equality of the transitive, reflexive, and symmetric closures of the given relational constraints. Related to the considered relations in this paper, this answers the cases for $r_{ab}$, $r_{com^{+}}$, and clearly also $r_{=}$. In contrast to that, we observe for other relations, namely $r_{rev}$, $r_{ssq}$ and $r_{*}$, that deciding equivalence in the non-erasing case cannot be characterized merely by such an equality property, showing a clear boundary to the already known trivial decidability of equivalence for $r_{=}$.
For the erasing case, most notably, we can show the existence of another relation besides $r_{|w|}$, namely $r_{\varepsilon_{=}}$, that allows for general decidability of the equivalence problem. On the other hand, for some of the considered relations, namely $_{rev}, rr_{*}$, and $r_{ssq}$, it can be shown that the equivalence problem must be at least as hard as the equivalence problem for $r_{=}$.
3.1 The Non-Erasing Case
Given any relational pattern $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, we denote by $r_{\alpha}]$ the reflexive, symmetric, and transitive closure on the set of binary related variables in $r_{\alpha}$. For example, if $(r_{=},x_{1},x_{2}),(r_{=},x_{2},x_{3})\in r_{\alpha}$, then we would have $(r_{=},x_{i},x_{j})\in[r_{\alpha}]$, for any combination $i,j\in\{1,2,3\}$. For the remainder of this section, we assume w.l.o.g. a normal form on the patterns such that all variables are always introduced inside a pattern with increasing indices, starting from 1 (i.e. $x_{1}$, $x_{2}$, $x_{3}$, and so on). The first results gives us a strong condition for relations $r$ that is reflexive on all letters of the alphabet and does not relate two distinct letters.
Lemma 1.
Assume $|\Sigma|\geq 2$ and let $R=\{r\}$, for a relation $r$ with the following properties: For all $\mathtt{a}\in\Sigma$, we have $(\mathtt{a},\mathtt{a})\in r$, and if $(\mathtt{a},\mathtt{b})\in r$, for two letters $\mathtt{a},\mathtt{b}\in\Sigma$, then $\mathtt{a}=\mathtt{b}$. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ be two patterns in normal form. If $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$ then $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$.
Proof.
First, we see that $|\alpha|=|\beta|$ must hold. Suppose otherwise and assume w.l.o.g. that $|\alpha|<|\beta|$. Take some $r_{\alpha}$-valid substitution $h$ that substitutes each variable by just a single letter. Then $|h(\alpha)|=|\alpha|$. In the NE-case, there can’t exist any $r_{\beta}$-valid substitution $h^{\prime}$ with $|h^{\prime}(\beta)|=|\alpha|<|\beta|$. A contradiction.
Now suppose $\alpha\neq\beta$. As both are regular patterns in normal form, that means there exists a first position $i\in[|\alpha|]$ such that w.l.o.g. $\alpha[i]=\mathtt{a}$, for some letter $\mathtt{a}\in\Sigma$, and $\beta[i]=x_{j}$, for some variable $x_{j}\in X$. Now let $h$ be a $r_{\beta}$-valid substitution that substitutes each variable by just a single letter and, in particular, set $h(x_{j})=\mathtt{b}$, for some letter $\mathtt{b}$ with $\mathtt{a}\neq\mathtt{b}$. As $|\alpha|=|\beta|$, any $r_{\alpha}$-valid substitution $h^{\prime}$ with $h^{\prime}(\alpha)0h(\beta)$ would have to match $h^{\prime}(\alpha[k])=h(\beta[k])$, for all $k\in[|\alpha|]$, and, in particular, would need to match $h^{\prime}(\alpha[i])=h^{\prime}(\mathtt{a})=\mathtt{a}\neq\mathtt{b}=h(x_{j})=h(\beta[i])$. A contradiction. That concludes $\alpha=\beta$.
Now suppose $r_{\alpha}]\neq[r_{\beta}]$. Then there exist two position $i,j\in[|\alpha|]$ such that $\alpha[i]=\beta[i]=x_{i^{\prime}}$, for some $i^{\prime}\in[|\alpha|]$, and $\alpha[j]=\beta[j]=x_{j^{\prime}}$, for some $j^{\prime}\in[|\alpha|]$, and w.l.o.g. $(r,x_{i^{\prime}},x_{j^{\prime}})\in[r_{\alpha}]$ but $(r,x_{i^{\prime}},x_{j^{\prime}})\notin[r_{\beta}]$. Now choose a substitution $h$ that substituted each variable just by a single letter such that $h(x_{i^{\prime}})=\mathtt{a}$ and $h(x_{j^{\prime}})=\mathtt{b}$, for two distinct letters $\mathtt{a},\mathtt{b}\in\Sigma$. As $r$ forces all blocks of symmetrically and transitively connected variables to be substituted by the same letter, in the case of length 1 substitutions, we know that $h$ cannot be $r_{\beta}$-valid. However, there exists such an $h$ that is $r_{\alpha}$-valid by choosing substitutions of the other variables accordingly. Hence, there exists such an $h(\alpha)\in L_{NE}(\alpha,r_{\beta})$, but we always have $h(\beta)\notin L_{NE}(\beta,r_{\beta})$. A contradiction. ∎
This result already confirms a very specific structure for all relations with the above stated properties. In particular, this covers the relations $r_{=}$, $r_{ab}$, $_{ssq}, rr$$_{rev}$, $r_{com^{*}}/r_{com^{*}}$ (equivalent for non-erasing pattern languages) and $r_{*}$. This property, however, is not exhaustive to characterize equivalence. So, adding to the previous property, if we now consider equivalence relations in general, we obtain the following result.
Lemma 2(✂0.A.1).
Let $R=\{r\}$ for an equivalence relation $r$. Let $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ be some relational pattern in normal form. We have $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\alpha,[r_{\alpha}])$.
Bringing the previous two kinds of relations together, we can immediately identify a set of relations for which equivalence of non-erasing languages can be characterized by equality of patterns and the corresponding closures.
Theorem 3.1(✂0.A.2).
Let $R=\{r\}$ for $r$ being an equivalence relation with the following properties: For all $\mathtt{a}\in\Sigma$, we have $(\mathtt{a},\mathtt{a})\in r$, and if $(\mathtt{a},\mathtt{b})\in r$, for two letters $\mathtt{a},\mathtt{b}\in\Sigma$, then $\mathtt{a}=\mathtt{b}$. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ be two patterns in normal form. $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$ if and only if $\alpha=\beta$ and $r_{\alpha}]=[r_{\beta}]$.
As the relations $r_{=}$, $r_{ab}$, and $r_{com^{+}}$ (also $r_{com^{*}}$ as it is the NE case) have the properties required in Theorem˜3.1, we can say the following.
Corollary 1(✂0.A.3).
@@ 3.1 The Non-Erasing Case
Let $R=\{r\}$ for some $r\in\{r_{=},r_{ab},r_{com^{+}}\}$. Given the relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, we can check in polynomial time whether $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\beta,r_{\beta})$.
Remark 1.
Notice that the relation $r_{\varepsilon_{=}}$ has no effect in the non-erasing case. In particular, for any relational pattern $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ with $R=\{r_{\varepsilon_{=}}\}$ we have $L_{NE}(\alpha,r_{\alpha})=L_{NE}(\alpha)$ for the regular pattern $\alpha$. Hence, equivalence between two such patterns is decidable in linear time in this case [3].
The other relations considered in Lemma˜1 that are not covered in Corollary˜1, namely $_{rev}, rr$$_{rev}$, $r_{*}$, are clearly not equivalence relations. Hence, decidability is not immediately clear in these cases. Here, blocks of related variables imply interesting structures on their own. Thus, they probably require a deeper investigation to answer (efficient) decidability in general (except the already known general decidability for $r_{rev}$ [30]). We conjecture they are all decidable.
3.2 The Erasing Case
In this part, our primary focus lies on the relation $r_{\varepsilon_{=}}$. First, note that $r_{\varepsilon_{=}}$ is reflexive, symmetric, and transitive. We extend the notation of the prior section by defining equivalence classes $\mathcal{C}_{\alpha}=\{C_{\alpha,1},\dots,C_{\alpha,k_{\alpha}}\}$ of variables $C_{\alpha,i}\subseteq\operatorname{var}(\alpha)$, $i\in[k_{\alpha}]$ that are in relation with one another. In the following, we prove that this relation results in decidability in the context of the erasing equivalence problem. For general patterns, we first present an polynomial space approach before giving linear-time and polynomial-time algorithms for patterns that satisfy specific properties.
Theorem 3.2(✂0.A.4).
Let $R=\{r_{\varepsilon_{=}}\}$ and $\Sigma$ be some alphabet. Given the relational patterns $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, checking $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$ is decidable in polynomial space.
Please consider Appendix 0.A.4 for a more detailed proof. Here, we provide a rough sketch. First, consider $(\alpha,r_{\alpha})$ and for each subset $S\subseteq\mathcal{C}_{\alpha}$, replace the variables $v\in\bigcup_{C_{\alpha,i}\in S}C_{\alpha,i}$ by any word in $\Sigma^{+}$ and any other variable $v\notin\bigcup_{C_{\alpha,i}\in S}C_{\alpha,i}$ by $\varepsilon$. We call the resulting language $L_{E}(\alpha,r_{\alpha},S)$. Using that, we obtain $L_{E}(\alpha,r_{\alpha})=\bigcup_{S\subseteq\mathcal{C}_{\alpha}}L_{E}(\alpha,r_{\alpha},S)$. Do the same for $\beta$. Note there exist at most $2^{k_{\alpha}}\leq 2^{|\alpha|}$ subsets $S$ for $\alpha$ and at most $2^{k_{\beta}}\leq 2^{|\beta|}$ subsets for $\beta$. For each $S\subseteq\mathcal{C}_{\alpha}$, we can take a sub-pattern $\alpha_{S}$ of $\alpha$ that only contains the variables that are not substituted by the empty word. As $\alpha_{S}$ is a regular pattern, it can be observed that $L_{E}(\alpha,r_{\alpha},S)=L_{NE}(\alpha_{s})$ and we can construct an NFA $A_{\alpha_{S}}$ with $L(A_{\alpha_{S}})=L_{NE}(\alpha_{S})=L_{NE}(\alpha,r_{\alpha},S)$. Doing that for each such $S$, combining all these NFAs into one big NFA $A_{\alpha}$ for $L_{E}(\alpha,r_{\alpha})$ and doing the same for $(\beta,r_{\beta})$ in an NFA $A_{\beta}$, we can compare the languages $L_{E}(\alpha,r_{\alpha})$ and $L_{E}(\beta,r_{\beta})$ by comparing $L(A_{\alpha})=L(A_{\beta})$. The latter is well known to be PSAPCE-complete (e.g., see [1]), concluding the result. Next, we introduce approaches with an improved running time for some pattern structures.
Lemma 3(✂0.A.5).
Let $R=\{r_{\varepsilon_{=}}\}$ and $\Sigma$ be some alphabet. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ be some relational patterns.
  1. 1.
    If $\alpha$ and $\beta$ are terminal-free, we can check in linear time whether $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$.
  2. 2.
    If $|\Sigma|=1$, we can check in linear time whether $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$.
  3. 3.
    If there exists $c\in\Sigma$ that is not a terminal in $\alpha$ and $\beta$, we can check in polynomial time whether $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$.
Finally, we consider some other relations in the context of the equivalence problem for erasing relational pattern languages. In particular, we can show that this problem must be at least as hard as in the case of $r_{=}$ for the relations $r_{*}$, $r_{ssq}$, and $r_{rev}$. While the first two follow trivially by simulating $r_{=}$, the last one requires a more technical construction (given in Appendix 0.A.6).
Proposition 1(✂0.A.6).
Let $R=\{r\}$ for $r\in\{r_{ssq},r_{*},r_{rev}\}$. Let $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$. Deciding $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$ is at least as hard as deciding the equivalence problem for erasing pattern languages.
@@ 4 The Inclusion Problem
In this section, we continue with the inclusion problem for relational pattern languages. We recall that the inclusion problem for non-erasing as well as erasing pattern languages over $r_{=}$ is generally undecidable for all alphabets $\Sigma$ of size $|\Sigma|\geq 2$, as shown and refined in  [23, 13, 5]. As we will see in this section, for several of the considered relations in $\mathcal{R}$, we also obtain that this problem is undecidable in the non-erasing or erasing cases. However, there also exist relations for which the problem is decidable.
First, consider the relations $r_{ssq}$ and $r_{*}$. We observe that we can simulate the exact behavior of $r_{=}$. Meaning, if we take any two variables $x,y\in\operatorname{var}(\alpha)$, for some relational pattern $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$, and any of the three relations above, we see that relating these variables in both directions effectively forces two variables to be substituted equally in $r_{\alpha}$-valid substitutions, i.e., if $(x,y),(y,x)\in r_{\alpha}$, then $h\in H_{r_{\alpha}}$ if and only if $h(x)=h(y)$. By the fact that this problem is undecidable for $r_{=}$, this results in the following.
Corollary 2.
[23, 13, 5] Let $\Sigma$ be any finite alphabet with $|\Sigma|\geq 2$. Let $r\in\{r_{=},r_{ssq},r_{*}\}$ and let $R=\{r\}$. Given $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$, it is undecidable to answer whether $L_{\mathcal{X}}(\alpha,r_{\alpha})\subseteq L_{\mathcal{X}}(\beta,r_{\beta})$, for $\mathcal{X}\in\{E,NE\}$.